OpenStudy (anonymous):

given a triangle with a=16, A=29 degrees, and B= 28 degreess, what is the lenth of c? a. 25.4 b. 22.4 c. 24.4 d. 23.4

OpenStudy (jdoe0001):

if \(\bf \measuredangle A=16^o\quad \measuredangle B=28^o\) what is the length of \(\bf \measuredangle C\quad ?\)

OpenStudy (jdoe0001):

when you find \(\bf \measuredangle C\), use the Law of Sines to get side "c"

OpenStudy (anonymous):

136

OpenStudy (jdoe0001):

hmmm wait... I... shoot... a =16... A = 29 =(

OpenStudy (jdoe0001):

anyhow your sum was correct either \(\bf \measuredangle A=29^o\quad \measuredangle B=28^o \quad \measuredangle C = 123^o\)

OpenStudy (jdoe0001):

using the Law of Sines \(\bf \cfrac{sin(123^o)}{c}=\cfrac{sin(29^o)}{16}\)

OpenStudy (anonymous):

is it B?

OpenStudy (jdoe0001):

... what did you get?

OpenStudy (anonymous):

honestly idk

OpenStudy (anonymous):

can you show me how to do it?!

OpenStudy (jdoe0001):

have you covered the Law of Sines yet?

OpenStudy (anonymous):

ya but i dont get it

OpenStudy (jdoe0001):

well... is very straighforward from \(\bf \cfrac{sin(123^o)}{c}=\cfrac{sin(29^o)}{16}\) just solve for "c"

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=It_bq5DzQ1Q